OpenStudy (anonymous):

One of the factors of 3p^5 – 12p^3 is: p^4 p + 2 p^2 + 4 4 I can't seem to solve this one, so yes help would be appreciated :)

OpenStudy (anonymous):

See, when you put p = 2, Then you are getting Zero..

OpenStudy (anonymous):

So, (p-2) is one of its factors.. Since it is not given in the choices, So by Long division method, divide given expression with (p-2)..

OpenStudy (anonymous):

1. What you wrote it an expression, not an equation. You solve equations, not expressions. 2. Factor out a 3p\(^3\). Then see if it factors further. (hint: it does indeed)

OpenStudy (anonymous):

K I get 3p^3(p^2-4p)

OpenStudy (anonymous):

@waterineyes , that's not an equation, you can't just plug in p = 2 like it's a function. Sorry m8

OpenStudy (anonymous):

(p^2-4)*

OpenStudy (anonymous):

You're correct so far @Loujoelou ;-)

OpenStudy (anonymous):

well almost

OpenStudy (anonymous):

3p^3(p^2-4)

OpenStudy (anonymous):

\[3p^3(p^2-4)\]

OpenStudy (anonymous):

Does x\(^2\) - a\(^2\) look familiar to you now? ;-)

OpenStudy (anonymous):

a = 2, a^2 = 4

OpenStudy (anonymous):

You can factor a bit more

OpenStudy (anonymous):

So, when you divide, You have now factors: \[(p-2)(3p^4 + 6p^3) = 3p^3.(p-2).(p+2) = 3p^3(p^2 - 4)\]

OpenStudy (anonymous):

So, P + 2 is the factor..

OpenStudy (anonymous):

@agentx5 I am doing it right..

OpenStudy (anonymous):

By equation you mean that you will find for x but I am not doing for that I am using that to make its factors...

OpenStudy (anonymous):

3p^3(p+2)(p-2) That's what I wrote, w/ (p+2) as answer but had to refresh OS cause froze on me.

OpenStudy (anonymous):

Tyvm @agentx5 & @waterineyes

OpenStudy (anonymous):

np m8

OpenStudy (anonymous):

Welcome dear..